# DeTPP: Leveraging Object Detection for Robust Long-Horizon Event Prediction

### v1

@@ Abstract
Forecasting future events over extended periods, known as long-horizon prediction, is a fundamental task in various domains, including retail, finance, healthcare, and social networks. Traditional methods, such as Marked Temporal Point Processes (MTPP), typically use autoregressive models to predict multiple future events. However, these models frequently encounter issues such as converging to constant or repetitive outputs, which significantly limits their effectiveness and applicability. To overcome these limitations, we propose DeTPP (Detection-based Temporal Point Processes), a novel approach inspired by object detection methods from computer vision. DeTPP utilizes a novel matching-based loss function that selectively focuses on reliably predictable events, enhancing both training robustness and inference diversity. Our method sets a new state-of-the-art in long-horizon event prediction, significantly outperforming existing MTPP and next-K approaches. The implementation of DeTPP is publicly available on GitHub  [^1] https://github.com/ivan-chai/hotpp-benchmark.
@@ 1 Introduction
The world is full of events. Internet activity, e-commerce transactions, retail operations, clinical visits, and numerous other aspects of our lives generate vast amounts of data in the form of timestamps and related information. These data points, when ordered by their timestamps, form what are known as Event Sequences (ESs). In the era of AI, it is crucial to develop methods capable of handling these complex data streams. Event sequences differ fundamentally from other data types. Unlike tabular data (Wang & Sun, 2022), ESs include timestamps and possess an inherent order. In contrast to time series data (Lim & Zohren, 2021), ESs are characterized by irregular time intervals and structured descriptions of each event. These differences necessitate the development of specialized models and evaluation practices.
@@
The primary task in the domain of ESs is sequence modeling, specifically predicting future event types and their occurrence times. Indeed, the ability to accurately forecast sequential events is vital for applications such as stock price prediction, personalized recommendation systems, and early disease detection. The simplified domain, represented as pairs of event types and times, is typically called Marked Temporal Point Processes (MTPP) (Rizoiu et al., 2017). Additionally, structured modeling of dependencies between different data fields (McDermott et al., 2024) can be considered an extension of MTPP.
@@
In practice, a common question arises: what events will occur, and when, within a specific time horizon? Forecasting multiple future events poses unique challenges distinct from traditional next-event prediction tasks. The conventional approach to this problem typically involves the autoregressive application of next-event prediction models. While these models have demonstrated effectiveness in predicting the immediate next event, their performance tends to decline as the prediction horizon extends (Karpukhin et al., 2024).
@@
In this paper, we show that autoregressive approach exhibit multiple problems when it comes to long-horizon prediction. In particular, an autoregressive model depends on its’ own errors and falls into constant or repetitive outputs. The other disadvantage is low inference parallelism due to the dependencies on latest predictions. We overcome these limitations by proposing DeTPP, a novel approach inspired by object detection techniques in computer vision. DeTPP predicts multiple future events in parallel. The novel matching loss, illustrated in Figure 1, skips events that are inherently hard to model and focuses on robust modeling of other events. This way, DeTPP achieves a new state-of-the-art in the long horizon prediction task, outperforming both autoregressive and next-K approaches.
@@
Figure 1: (a) A typical next-item or next-K loss compares events at corresponding positions, often resulting in incorrect matches. (b) The proposed matching loss calculates the loss function between the closest events, leading to a more robust and balanced error measure. (c) The proposed DeTPP method enhances prediction diversity. The illustration shows example predicted labels from the Amazon dataset for the autoregressive MAE-CE method, its Next-K extension, and the proposed DeTPP method. Each label type is depicted using a distinct shape and color combination.
2 Background
@@ 2.1 Marked Temporal Point Processes
Marked Temporal Point Process (MTPP) is a stochastic process that consists of a sequence of time-event pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ denote the times of events, and $l_{i}\in\{1,\dots,L\}$ are the corresponding event type labels. Traditional MTPP models primarily focus on predicting the next event in the sequence. A straightforward approach is to independently predict the time and type of the next event, while more sophisticated methods model the temporal dynamics of each event type separately. These models often rely on Temporal Point Processes (Rizoiu et al., 2017), which describe the stochastic generation of event times.
@@ 2.2 Models
Temporal Point Processes. Traditional Marked Temporal Point Process (MTPP) models, such as Poisson and Hawkes processes (Rizoiu et al., 2017), rely on strong assumptions about the underlying generative processes. Recent advancements have shifted towards more flexible and expressive models by leveraging neural architectures. These include classical Recurrent Neural Networks (RNNs) (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019; Shchur et al., 2019), as well as more advanced architectures like transformers (Zhang et al., 2020; Zuo et al., 2020; Wang & Xiao, 2022; Yang et al., 2022). Additionally, continuous-time models such as Neural Hawkes Processes (Mei & Eisner, 2017), ODE-RNN (Rubanova et al., 2019), and their variants (Jia & Benson, 2019; De Brouwer et al., 2019; Kidger et al., 2020; Song et al., 2024; Kuleshov et al., 2024) have been developed to better capture the dynamics of event sequences. Moreover, some researchers have adapted generative models, including denoising diffusion and Generative Adversarial Networks (GANs), for use in TPPs (Lin et al., 2022).
@@
Long-horizon models. The problem of long-horizon prediction has been specifically addressed by HYPRO (Xue et al., 2022), which introduces a technique for selecting the best candidate from a set of generated sequences. HYPRO functions as a meta-algorithm that can enhance the performance of nearly any sequence prediction model. However, HYPRO’s approach requires multiple generation runs for each prediction, which significantly reduces training and inference speed.
@@ 2.3 Evaluation
MTPP models are typically evaluated based on their accuracy in predicting the next event (Xue et al., 2023). Time and type predictions are often assessed separately, with type prediction accuracy measured by error rates and time prediction accuracy evaluated using metrics like Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE). Recent advancements have introduced metrics such as Optimal Transport Distance (OTD) (Mei et al., 2019) and Temporal mAP (T-mAP) (Karpukhin et al., 2024), which are designed to evaluate long-horizon predictions by comparing predicted sequences to ground truth sequences within a specified horizon. In this work, we use all mentioned metrics to evaluate the performance of our proposed method.
@@ 3 Limitations of Autoregressive Inference
Previous studies have identified several challenges associated with autoregressive models for long-horizon predictions (Karpukhin et al., 2024). Notably, these models often exhibit reduced prediction uncertainty over extended horizons, even though the task becomes increasingly difficult. As illustrated in Figure 1.c, the predicted label sequences often exhibit a tendency towards constant or repetitive outputs. This behavior likely stems from the model’s reliance on its own predictions as input for subsequent predictions, which can amplify errors and lead to the repetition of events. Interestingly, even the Next-K approach, which avoids autoregressive dependencies, suffers from similar repetitive patterns. This may be attributed to each head in the Next-K model predicting the entire distribution of labels, leading to a bias toward the most frequent classes during inference.
In contrast, our proposed DeTPP approach exhibits greater diversity in its predictions and, as we will demonstrate, achieves superior performance in long-horizon prediction tasks.
@@ 4 Event Detection with DeTPP
In this section, we introduce DeTPP, a novel approach to MTPP modeling that leverages concepts from object detection (Carion et al., 2020). DeTPP is designed to predict $\mathrm{K}$ future events within a specified time horizon $\mathrm{H}$, where $\mathrm{K}$ and $\mathrm{H}$ are hyperparameters of the model. The generative process of DeTPP is illustrated in Figure 2. Below, we provide a detailed description of the model, along with its training and inference procedures.
@@ 4.1 Probabilistic Event Model
As depicted in Figure 2, the probability $\hat{o}$ of an event occurring is modeled using a neural network head with a sigmoid activation function. A separate head with softmax activation models the distribution $\hat{p}(l)$ of event labels. For the time shift, we employ an approach similar to Mixture Density Networks (MDNs) (Bishop, 1994), modeling the time shift as a Laplace distribution with a unit scale parameter:
@@
where $\hat{t}$ is the model’s predicted time shift. This formulation provides a probabilistic interpretation of the Mean Absolute Error (MAE) loss function. By combining these predicted probabilities, we can estimate the likelihood of a future event given the model’s output:
@@
$\log\mathrm{P}(\emptyset)=\log(1-\hat{o}).$ (3)
This probabilistic framework provides a rigorous basis for evaluating the likelihood of different event outcomes.
Figure 2: DeTPP predicts K future events. Each prediction includes presence probability $\hat{o}$, time $\hat{t}$, and labels distribution $\hat{p}(l)$. During training a special matching loss aligns predictions with ground truth sequence and evaluates its likelihood.
4.2 Horizon Modeling approach
@@
where $\mathcal{A}$ is the set of all possible alignments between the ground truth and predicted sequences, and $\sigma$ represents a specific alignment. The optimal matching is computed using the Hungarian algorithm, which efficiently solves this assignment problem.
@@
where $y=(t,l)$ is a ground truth event, $\hat{t}$ is the predicted timestamp, and $\hat{p}(l)$ is the predicted probability of the correct label.
The binary cross-entropy loss $\mathcal{L}_{\mathrm{BCE}}$ is used to train the model to predict the presence probability of events:
@@
By minimizing the overall loss $\mathcal{L}(y,\hat{y})$, DeTPP is trained to accurately predict the parameters of the ground truth sequence, adapting to sequences of varying lengths up to the maximum of $\mathrm{K}$ events.
4.3 Inference
Inference with DeTPP involves two steps: filtering and sorting. Initially, predicted events are filtered based on their presence probabilities $\hat{o}_{i}$. Only events with a presence probability above a certain threshold are retained. The remaining events are then sorted according to their predicted timestamps, forming the final output sequence.
@@ 4.3 Inference
However, in practice, an additional calibration step is necessary. Without calibration, the model tends to predict few events due to conservative probability estimates. The goal of calibration is to determine optimal thresholds for each $\hat{o}_{i}$, aligning them with the prior probability of the event $i$ being matched. This calibration is performed on-the-fly during training by tracking matching frequencies and computing the corresponding quantiles using a streaming algorithm.
Table 1: Evaluation results. The best result is shown in bold. Mean and STD values of 5 runs with different random seeds are reported.
Method Mean length Labels entropy Time delta diff. entropy Next-item Acc., % Next-item MAE OTD Val / Test T-mAP, % Val / Test Retweet MAE-CE 15.0 0.45 3.38 59.95$\pm$ 0.03 24.31$\pm$ 0.04 173.3$\pm$ 4.1 / 172.7$\pm$ 4.4 33.49$\pm$ 5.07 / 31.75$\pm$ 4.95 RMTPP 13.3 0.51 3.63 60.07$\pm$ 0.10 25.52$\pm$ 0.17 167.7$\pm$ 3.4 / 166.6$\pm$ 3.4 48.22$\pm$ 2.34 / 45.41$\pm$ 2.49 NHP 13.3 0.45 3.67 60.09$\pm$ 0.03 25.50$\pm$ 0.11 167.0$\pm$ 1.5 / 165.7$\pm$ 1.6 49.57$\pm$ 0.95 / 46.91$\pm$ 1.01 ODE 13.5 0.49 3.62 59.94$\pm$ 0.07 25.47$\pm$ 0.03 166.6$\pm$ 0.4 / 165.3$\pm$ 0.5 48.96$\pm$ 0.73 / 45.96$\pm$ 0.94 MAE-CE-K 14.8 0.53 3.14 59.55$\pm$ 0.26 25.99$\pm$ 0.81 169.1$\pm$ 2.2 / 168.5$\pm$ 2.6 36.64$\pm$ 4.95 / 34.27$\pm$ 5.15 DeTPP 23.9 0.76 2.64 55.83$\pm$ 1.23 28.76$\pm$ 0.10 147.2$\pm$ 0.3 / 145.2$\pm$ 0.3 59.31$\pm$ 0.34 / 55.78$\pm$ 0.47 Amazon MAE-CE 13.2 1.06 -1.76 35.73$\pm$ 0.12 0.242$\pm$ 0.003 6.62$\pm$ 0.05 / 6.55$\pm$ 0.04 21.99$\pm$ 0.38 / 22.60$\pm$ 0.35 RMTPP 15.3 1.19 -0.64 35.75$\pm$ 0.05 0.294$\pm$ 0.001 6.67$\pm$ 0.02 / 6.62$\pm$ 0.02 20.06$\pm$ 0.29 / 20.42$\pm$ 0.30 NHP 9.9 1.56 -0.92 11.00$\pm$ 3.47 0.449$\pm$ 0.014 9.05$\pm$ 0.30 / 9.04$\pm$ 0.34 25.05$\pm$ 0.46 / 24.91$\pm$ 0.60 ODE 9.5 1.84 -0.81 7.61$\pm$ 0.99 0.492$\pm$ 0.018 9.47$\pm$ 0.11 / 9.48$\pm$ 0.08 22.50$\pm$ 0.63 / 21.88$\pm$ 0.77 MAE-CE-K 13.3 1.06 -2.24 35.11$\pm$ 0.08 0.246$\pm$ 0.000 6.74$\pm$ 0.01 / 6.70$\pm$ 0.01 21.84$\pm$ 0.05 / 22.35$\pm$ 0.07 DeTPP 18.6 1.72 0.11 31.94$\pm$ 1.33 0.264$\pm$ 0.012 6.33$\pm$ 0.04 / 6.30$\pm$ 0.04 37.80$\pm$ 0.15 / 37.97$\pm$ 0.14 StackOverflow MAE-CE 15.7 0.59 -0.31 45.41$\pm$ 0.11 0.641$\pm$ 0.002 13.58$\pm$ 0.07 / 13.65$\pm$ 0.05 8.85$\pm$ 0.66 / 8.46$\pm$ 0.39 RMTPP 12.5 0.56 -0.03 45.43$\pm$ 0.13 0.700$\pm$ 0.007 12.97$\pm$ 0.02 / 13.18$\pm$ 0.05 13.39$\pm$ 0.23 / 12.99$\pm$ 0.11 NHP 11.9 0.44 0.02 44.54$\pm$ 0.05 0.716$\pm$ 0.004 13.02$\pm$ 0.02 / 13.25$\pm$ 0.02 12.62$\pm$ 0.55 / 11.93$\pm$ 0.32 ODE 12.1 0.38 -0.02 44.37$\pm$ 0.07 0.713$\pm$ 0.003 13.04$\pm$ 0.02 / 13.28$\pm$ 0.03 11.25$\pm$ 0.35 / 10.50$\pm$ 0.28 MAE-CE-K 15.3 0.47 -0.69 44.85$\pm$ 0.24 0.644$\pm$ 0.003 13.41$\pm$ 0.08 / 13.52$\pm$ 0.06 12.14$\pm$ 0.99 / 11.17$\pm$ 0.79 DeTPP 12.1 1.16 0.50 39.98$\pm$ 1.95 0.783$\pm$ 0.025 12.99$\pm$ 0.05 / 13.14$\pm$ 0.05 18.66$\pm$ 0.33 / 17.80$\pm$ 0.30

### v2

# DeTPP: Leveraging Object Detection for Robust Long-Horizon Events Forecasting

@@ Abstract
Long-horizon event forecasting is critical across various domains, including retail, finance, healthcare, and social networks. Traditional methods, such as Marked Temporal Point Processes (MTPP), often rely on autoregressive models to predict multiple future events. However, these models frequently suffer from issues like converging to constant or repetitive outputs, which limits their effectiveness and general applicability. To address these challenges, we introduce DeTPP (Detection-based Temporal Point Processes), a novel approach inspired by object detection techniques from computer vision. DeTPP employs a unique matching-based loss function that selectively prioritizes reliably predictable events, improving the accuracy and diversity of predictions during inference. Our method establishes a new state-of-the-art in long-horizon event forecasting, achieving up to a 77% relative improvement over existing MTPP and next-K methods. The proposed hybrid approach enhances the accuracy of next event prediction by up to 2.7% on a large transactional dataset. Notably, DeTPP is also among the fastest methods for inference. The implementation of DeTPP is publicly available on GitHub [^1] https://github.com/ivan-chai/hotpp-benchmark.
@@ 1 Introduction
Internet activity, e-commerce transactions, retail operations, clinical visits, and many other aspects of our lives generate vast amounts of data in the form of timestamps and related information. When ordered by their timestamps, these data points form event sequences, and it is crucial to develop methods capable of handling these complex data streams. Event sequences differ fundamentally from other data types. Unlike tabular data (Wang & Sun, 2022), events include timestamps and have an inherent order. In contrast to time series data (Lim & Zohren, 2021), event sequences are characterized by irregular time intervals and additional attributes. These distinctions necessitate the development of specialized models.
@@
The primary task in the domain of event sequences is predicting future event types and their occurrence times. Indeed, the ability to accurately forecast sequential events is vital for applications such as stock price prediction, personalized recommendation systems, and early disease detection. The simplified domain, represented as pairs of event types and times, is typically called Marked Temporal Point Processes (MTPP) (Rizoiu et al., 2017). Additionally, structured modeling of dependencies between different data fields (McDermott et al., 2024) can be considered an extension of MTPP.
@@
Practical applications often require predicting multiple future events within a specified time horizon. This task presents unique challenges that differ from traditional next-event prediction. The conventional approach typically relies on autoregressive models, which predict the next event step by step. While these models are effective for immediate next-event forecasting, their performance tends to deteriorate as the prediction horizon extends (Karpukhin et al., 2024).
@@
This paper demonstrates that the autoregressive approach faces several issues in long-horizon prediction. In particular, autoregressive models accumulate errors over time, leading to constant or repetitive outputs. Additionally, their inference parallelism is limited due to dependencies on the latest predictions. To address these limitations, we propose DeTPP, a novel approach inspired by object detection techniques in computer vision. DeTPP predicts multiple future events in parallel. The novel matching loss, illustrated in Fig. 1, skips events that are inherently hard to model and focuses on robust modeling of other events. As a result, DeTPP sets a new state-of-the-art in long-horizon prediction, outperforming both autoregressive and next-K approaches. We also introduce a simple extension that combines ideas from DeTPP with traditional methods, improving next-event prediction quality, particularly on the large Transactions dataset.
@@
Figure 1: (a) A typical next-item or next-K loss compares events at corresponding positions, often resulting in incorrect matches. (b) The proposed matching loss calculates the loss function between the closest events, leading to a more robust and balanced error measure. (c) The proposed DeTPP method enhances prediction diversity. The illustration shows example predicted labels from the Amazon dataset for the autoregressive IFTPP method, its next-K extension, and the proposed DeTPP method. Each label type is depicted using a distinct shape and color combination.
2 Related Works
@@ 2.1 Marked Temporal Point Processes
MTPP is a stochastic process that consists of a sequence of time-event pairs $(t_{1},l_{1}),(t_{2},l_{2}),\dots$, where $t_{1}<t_{2}<\dots$ denote the times of events, and $l_{i}\in\{1,\dots,L\}$ are the corresponding event type labels (Rizoiu et al., 2017). Traditional MTPP models primarily focus on predicting the next event in the sequence. A straightforward approach is to independently predict the time and type of the next event, while more sophisticated methods model the temporal dynamics of each event type separately. These models often rely on Temporal Point Processes, which describe the stochastic generation of event times.
@@ 2.2 Models
Marked Temporal Point Processes. Traditional MTPP models, such as Poisson and Hawkes processes (Rizoiu et al., 2017), rely on strong assumptions about the underlying generative processes. Recent advancements have shifted towards more flexible and expressive models that leverage neural architectures. These include classical Recurrent Neural Networks (RNNs) (Du et al., 2016; Xiao et al., 2017; Omi et al., 2019; Shchur et al., 2019), as well as more advanced architectures like transformers (Zhang et al., 2020; Zuo et al., 2020; Wang & Xiao, 2022; Yang et al., 2022). Additionally, continuous-time models such as Neural Hawkes Processes (Mei & Eisner, 2017), ODE-RNN (Rubanova et al., 2019), and their variants (Jia & Benson, 2019; De Brouwer et al., 2019; Kidger et al., 2020; Song et al., 2024; Kuleshov et al., 2024) have been developed to capture the dynamics of event sequences better. Moreover, some researchers have adapted generative models, including denoising diffusion and Generative Adversarial Networks (GANs), for use in TPPs (Lin et al., 2022).
@@
Long-horizon models. The problem of long-horizon prediction has been addressed explicitly by HYPRO (Xue et al., 2022), which introduces a technique for selecting the best candidate from a set of generated sequences. HYPRO functions as a meta-algorithm that can enhance the performance of nearly any sequence prediction model. However, HYPRO’s approach requires multiple generation runs for each prediction, significantly reducing training and inference speed.
@@ 2.3 Evaluation
MTPP models are typically evaluated based on their accuracy in predicting the next event (Xue et al., 2023). Time and type predictions are often assessed separately, with type prediction accuracy measured by error rates and time prediction accuracy evaluated using metrics like Mean Absolute Error (MAE) or Root Mean Squared Error (RMSE). Recent advancements have introduced metrics such as Optimal Transport Distance (OTD) (Mei et al., 2019) and Temporal mAP (T-mAP) (Karpukhin et al., 2024), which are designed to evaluate long-horizon predictions by comparing predicted sequences to ground truth sequences within a specified horizon. In this work, we use all mentioned metrics to assess the performance of our proposed method. We also measure the diversity of predictions by using the predicted labels distribution entropy.
@@ 3 Limitations of Autoregressive Inference
Previous studies have identified several challenges associated with autoregressive models for long-horizon predictions (Karpukhin et al., 2024). These models often exhibit reduced prediction uncertainty over extended horizons, even though the task becomes increasingly difficult. As illustrated in Fig. 1.c, the predicted label sequences often have constant or repetitive outputs. This behavior likely stems from the model’s reliance on its predictions as input for subsequent predictions, which can amplify errors and lead to repeated events. Interestingly, even the next-K approach, which avoids autoregressive dependencies, suffers from similar repetitive patterns. This may be attributed to each head in the next-K model predicting the entire distribution of labels, leading to a bias toward the most frequent classes during inference.
In contrast, our proposed DeTPP approach exhibits greater diversity in its predictions and, as we will demonstrate, achieves superior performance in long-horizon prediction tasks. Let us discuss the details of the proposed method in the next Section.
@@ 4 Event Detection with DeTPP
This section introduces DeTPP, a novel approach to MTPP modeling that incorporates concepts from object detection (Carion et al., 2020). DeTPP is designed to predict $\mathrm{K}$ future event candidates within a specified time horizon $\mathrm{H}$, where $\mathrm{K}$ is larger than the typical sequence length. During training, the model aligns its predictions with ground truth and computes pairwise losses, as illustrated in Fig. 2. At inference time, the model retains only candidates with high prediction scores. Below, we provide a detailed description of the model, as well as its training and inference procedures.
@@ 4.1 Probabilistic Event Model
As depicted in Fig. 2, the probability $\hat{o}$ of an event occurring is modeled using a neural network with a sigmoid activation function. A separate head with softmax activation models the distribution $\hat{p}(l)$ of event labels. For the time shift, we employ an approach similar to Mixture Density Networks (MDNs) (Bishop, 1994) and intensity-free TPP (Shchur et al., 2019), modeling the time shift as a Laplace distribution with a unit scale parameter:
@@
where $\hat{t}$ is the predicted time shift. This formulation offers a probabilistic interpretation of the MAE loss function. By combining the predicted probabilities, we can estimate the likelihood of a future event given the output of the model:
@@
$\log\mathrm{P}(\emptyset)=\log(1-\hat{o})+C_{\emptyset},$ (3)
where $C_{\emptyset}$ is a constant independent of the model’s output, representing the probability associated with a reserved ”unknown” time and label values. To compute the loss function, we omit $C_{\emptyset}$ since it does not influence the gradient during training.
This proposed probabilistic framework offers a rigorous basis for evaluating the likelihood of ground truth event sequences.
Figure 2: DeTPP predicts K future events. Each prediction includes presence probability $\hat{o}$, time $\hat{t}$, and labels distribution $\hat{p}(l)$. During training, a special matching loss aligns predictions with the ground truth sequence and evaluates its likelihood.
4.2 Horizon Matching Loss
@@
where $\mathcal{A}$ is the set of all possible alignments between the ground truth and predicted sequences and $\sigma$ represents a specific alignment. The optimal matching is computed using the Hungarian algorithm.
@@
where $y=(t,l)$ is a ground truth event, $\hat{t}$ is the predicted timestamp, and $\hat{p}(l)$ is the predicted probability of the correct label. The binary cross-entropy loss $\mathcal{L}_{\mathrm{BCE}}$ is used to train the model to predict the presence probability of events:
@@
By minimizing the overall loss $\mathcal{L}(y,\hat{y})$, DeTPP is trained to accurately predict the parameters of the ground truth sequence, adapting to sequences of varying lengths, up to a maximum of $\mathrm{K}$ events. Unlike object detection models such as DeTR Carion et al. (2020), DeTPP employs the same loss function for both the matching process and model training. Specifically, DeTPP integrates the alignment loss $\mathcal{L}_{\mathrm{BCE}}$ into the matching cost. As demonstrated in our experiments, this approach enhances the model’s performance, especially on small datasets.
4.3 Conditional Head Architecture
Figure 3: The DeTPP conditional prediction head.
Implementing a separate feed-forward network for each prediction head results in many parameters and increases the risk of overfitting. A common strategy to reduce the parameter count is applying a transformer decoder network to a set of $K$ queries, where the decoder shares the same projection matrices for all input vectors, lowering the total number of weights. However, since our model generates the entire set of predictions from a single embedding vector, we can simplify this approach by omitting the cross-attention layer.
In DeTPP, we implement a conditional feed-forward network, as shown in Fig. 3. The model takes two inputs: a query and a context vector. It then applies a feed-forward network to the concatenated input vectors. This way, the trainable query vector encodes the necessary information about each output head, allowing the same feed-forward network to generate all $K$ outputs. Using a single network instead of $K$ distinct ones significantly reduces the overall number of parameters, speeds up convergence, and improves prediction quality, as demonstrated in our ablation studies.
4.4 Calibration and Inference
@@
Inference with DeTPP involves two steps: filtering and sorting. Initially, predicted events are filtered based on their presence probabilities $\hat{o}_{i}$. The remaining events are sorted according to their predicted timestamps, forming the final output sequence. However, in practice, an additional calibration step is necessary. The model tends to predict a few events without calibration due to conservative probability estimates. Calibration aims to determine optimal prediction thresholds for all $\hat{o}_{i}$, aligning prediction rates with matching probabilities. This calibration is performed on the fly by tracking matching frequencies and computing the corresponding quantiles using a streaming algorithm. The proposed calibration algorithm is outlined in Appendix A.
4.5 Hyperparameter Selection
DeTPP relies on two key hyperparameters: the maximum number of predictions $K$ and the horizon $H$. We set the horizon $H$ to align with that of the T-mAP metric, ensuring consistency in evaluation. The selected value is also sufficient to include the necessary number of events for calculating the OTD metric. The selection of $K$ typically requires tuning for each dataset. It controls the maximum number of predictions, and we recommend setting $K$ to approximately four times the average sequence length. In our experiments, $K$ values ranged between 32 and 64, depending on the dataset’s characteristics.
Additionally, we found it beneficial to adjust the weight of each loss function during alignment to better accommodate the number of model outputs, dataset classes, and the average time step. In practice, the optimal weight for $\mathcal{L}_{\mathrm{BCE}}$ is typically around eight times larger than the weights for the label and timestamp losses. For datasets with larger time steps, such as Retweet, the MAE loss weight should be reduced accordingly.
The exact hyperparameter values used in our experiments are provided in Appendix B.